OpenStudy (anonymous):

Kim surveyed 80 students at her school and found that 0.65 of the students she surveyed said their favorite class is math. Another 15% of the students she surveyed reported that their favorite class is science. How many more students in the survey prefer math over science?

OpenStudy (freckles):

well 65% of 80 preferred math what is .65(80) and 15% of 80 preferred science what is .15(80)

OpenStudy (anonymous):

12

OpenStudy (freckles):

then find the difference

OpenStudy (freckles):

.15(80)=12 .65(80)=?

OpenStudy (anonymous):

52

OpenStudy (freckles):

that means 12 people preferred science (since .15*80=12) now we want to know how many of the 80 preferred math (this is where you need to calculate .65*80

OpenStudy (freckles):

and yes

OpenStudy (freckles):

so you have 12 people prefer science and 52 prefer math

OpenStudy (anonymous):

.65(80) is 52

OpenStudy (anonymous):

so now what do i do?

OpenStudy (freckles):

find the difference

OpenStudy (anonymous):

so 52-12=30?

OpenStudy (freckles):

5-1=4 I believe so 52-12=40 so 40 people preferred math over science

OpenStudy (anonymous):

THANK YOU SO MUCH 4 THE HELP

OpenStudy (freckles):

np